OpenStudy (gabylovesyou):

While simplifying some math work, Peter wrote on his paper that x^3 • x^3 • x^3 • x^3 equaled x^3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x^3 + x^3 + x^3 + x^3?

OpenStudy (minecrafter64):

no

OpenStudy (diamondboy):

peter didn't simplify his work completely and peters work is not the same as x^3+x^3+x^3+x^3

OpenStudy (minecrafter64):

it wouldnt be the same

OpenStudy (diamondboy):

this is d simplified answer x^12

OpenStudy (minecrafter64):

true dat

OpenStudy (minecrafter64):

x^12 is the answer all in agreement.

OpenStudy (nurali):

https://prezi.com/3dsk70wt_tnv/202/

OpenStudy (mathstudent55):

The first question is correct, only it has to be written correctly. \(\LARGE x^3 \cdot x^3 \cdot x^3 \cdot x^3 = x^{3+3+3+3} \) The above line is correct. That was your first question.

OpenStudy (mathstudent55):

The above line was correct, but it was not a final answer. A final answer would be (below, in red): \(\LARGE x^3 \cdot x^3 \cdot x^3 \cdot x^3 = x^{3+3+3+3} = \color{red}{x^{12}}\)

OpenStudy (mathstudent55):

The second question is: What is \(\LARGE x^3 + x^3 + x^3 + x^3 = \)

OpenStudy (mathstudent55):

This is a completely different question. Here you need to add terms together. You can only add like terms together. Like terms have the same variables and the same exponents. You see that you have 4 terms. All of them are the same, x^3. That means all terms are like terms. You can add them together. \(\LARGE x^3 + x^3 + x^3 + x^3 = 1x^3 + 1x^3 + 1x^3 + 1x^3 =\) \(\LARGE =x^3(1 + 1 + 1 + 1) = x^3 \cdot 4 = 4x^3\) The first question's answer was \(\LARGE x^{12}\) Now you see that the answers are very different.